# Splittings in a complete local ring and decomposition its group of units

### v3

@@ Abstract.
Let $(R,M,k)$ be a complete local ring (not necessarily Noetherian). Then we prove that the natural surjective ring map $R\rightarrow k$ admits a splitting if and only if $\operatorname{char}(R)=\operatorname{char}(k)$. In addition, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then we prove that the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
We also show with an example that the above theorem is in its full generality. As an application of this theorem, we show that the following short exact sequence of Abelian groups:
@@ 1. Introduction
Then, after exploring several stages of the level of generalization, first for Artinian local rings, then for local rings whose Jacobson radical are nilpotent, finally as one of the main results of this investigation, we succeeded to show that the same splitting holds for any complete local ring. However, the proof of this general case is technical, and uses the tools and methods of commutative algebra, the theory of fields, and homological algebra. In fact, we prove a more general result:
@@
In proving this theorem, as is clear from its formulation, the characteristic of the residue field $k=R/M$ plays an important role and hence it is necessary to examine it carefully. We know that in any local ring $(R,M,k)$, if $\operatorname{char}(R)=\operatorname{char}(k)$ then there are two cases: $\operatorname{char}(k)=0$ or $\operatorname{char}(k)=p$ for some prime number $p$. However, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then the characteristic of $k$ is always a prime number $p$. To prove the above theorem, in the first case, i.e., when $\operatorname{char}(R)=\operatorname{char}(k)$, one of the main ingredients that applied is the following important result:
@@
To prove Theorem 1.1, in the second case, i.e., when $\operatorname{char}(R)\neq\operatorname{char}(k)$, one of the main tools which used is a key result in homological algebra on the vanishing of Ext groups:
@@ 2. Finite case-the initial motivation
Also recall that if $f:A\rightarrow B$ is a morphism of rings, then it can be seen that $\operatorname{char}(B)$ divides $\operatorname{char}(A)$. In particular, if we have morphisms of rings $A\rightarrow B$ and $B\rightarrow A$, then $\operatorname{char}(A)=\operatorname{char}(B)$. If we have a ring map $A\rightarrow B$ with $\operatorname{char}(B)=0$, then $\operatorname{char}(A)=0$. However, if $\operatorname{char}(A)=p^{n}$ for some prime number $p$, then $\operatorname{char}(B)=p^{d}$ with $d\leqslant n$.
@@
Recall that by a complete local ring we mean a local ring $(R,M)$ such that it is complete with respect to the $M$-adic topology (i.e., $R$ is Hausdorff and every Cauchy sequence is convergent in $R$), or equivalently, the natural ring map $R\rightarrow\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ given by $r\mapsto(r+M^{n})_{n\geqslant 1}$ is an isomorphism. Note that in this definition $R$ is not assumed to be Noetherian (contrary to the classical literature, see e.g. [1], where complete local rings were also assumed to be Noetherian).
@@ 3. The general case
Using the above isomorphism of Abelian groups we can make $G$ into an $R/I$-module (and then this map becomes an isomorphism of $R/I$-modules). Now the statement about the vanishing is clear, because if $M$ is a module over a ring $R$ then $M$ is vanished by $\operatorname{char}(R)$. ∎
@@
The above corollary is quite sufficient for our purpose. However, we should add that this result holds more generally: Let $R$ be a regular Noetherian ring of Krull dimension $n$ and $M,N$ be modules over $R$. If there is an element $r\in R$ such that $r$ is regular on $M$ and $rN=0$, then $\operatorname{Ext}_{R}^{n}(M,N)=0$. In fact, this is proved by the same argument as the above by applying a theorem of Serre [9] which asserts that if $R$ is regular Noetherian of Krull dimension $n$, then for any two $R$-modules $M,N$ one has $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>n$. The above corollary also holds over a Dedekind domain as well.
@@
Now let $p$ be a prime number. Consider the local ring $\mathbb{Z}_{P}$ where $P=p\mathbb{Z}$. By the above, $\mathbb{Z}_{P}/P\mathbb{Z}_{P}\simeq\mathbb{F}_{p}=\mathbb{Z}/p$. Then the natural surjective map between the groups of units $\mathbb{Z}_{P}^{\ast}\rightarrow\mathbb{F}_{p}^{\ast}$ does not admit a splitting for $p\geqslant 5$, because if this map splits then $\mathbb{F}_{p}^{\ast}$ can be viewed as a subgroup of $\mathbb{Z}_{P}^{\ast}$, but $\mathbb{Z}_{P}$ does not contain roots of unit of order p1p-1.
But for $p=2,3$ the map $\mathbb{Z}_{P}^{\ast}\rightarrow\mathbb{F}_{p}^{\ast}$ admits a splitting. We know that for any prime number $p$, the local ring $\mathbb{Z}_{P}$ is not complete. This shows that the above short exact sequence may be split for some incomplete local rings (in other words, the converse of Corollary 3.16 is not true).
@@
First of all, note that if a ring $R$ is complete with respect to the $I$-adic topology for some ideal $I$ of $R$, then it can be seen that $I$ is contained in the Jacobson radical of $R$. In particular, if a ring $R$ is complete with respect to the $M$-adic topology for some maximal ideal $M$ of $R$, then $R$ is a (complete) local ring, and for complete local rings we proved that the above splittings already exist. Hence, to find the above examples, the ideal $I$ must be a non-maximal ideal.
@@
For the second type of example, for a prime number $p$, consider the ring of $p$-adic integers $R=\mathbb{Z}_{p}=\lim\limits_{\overleftarrow{n\geqslant 1}}\mathbb{Z}/p^{n}$ which is a complete local ring of mixed characteristic, i.e., $\operatorname{char}(\mathbb{Z}_{p})=0$ but the reside field of $\mathbb{Z}_{p}$ is $\mathbb{F}_{p}$, and take the ideal $I=(p^{2})$ of $R$ where by $p^{2}$ we mean the element $(p^{2}+p^{n}\mathbb{Z})_{n\geqslant 1}$ of $R$. By Remark 2.5, we have $R/I\simeq\mathbb{Z}/p^{2}$. Then the group of units of $R/I$ is of order p(p1)p(p-1) and by Cauchy theorem has an element of order $p$ (it can be also directly seen that $1+p$ is an element of $(\mathbb{Z}/p^{2})^{\ast}$ of order $p$), but $(\mathbb{Z}_{p})^{\ast}$ has no nontrivial $p$-torsion for $p\geqslant 3$, so there is no splitting in this case.
Acknowledgments. The author would like to give sincere thanks to Professors Bhargav Bhatt, Brian Conrad, Pierre Deligne, and Melvin Hochster who generously shared their very valuable and excellent ideas with me during the writing of this article, and for their important role in proving the main results of this article.

### v4

# Splitting in a complete local ring and decomposition its group of units

@@ Abstract.
Let $(R,M,k)$ be a complete local ring (not necessarily Noetherian). Then we reprove by a new method that the natural surjective ring map $R\rightarrow k$ admits a splitting if and only if $\operatorname{char}(R)=\operatorname{char}(k)$. In our proof there is no need for the existence of the coefficient field for equi-characteristic complete local rings, whose existence is the hardest part of the known proof.
However, the main result of this article is that in the unequal characteristic case $\operatorname{char}(R)\neq\operatorname{char}(k)$, we prove that the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
As an application of the above theorem, we show that for any complete local ring $(R,M,k)$ the following short exact sequence of Abelian groups:
@@ 1. Introduction
Then, after exploring several stages of the level of generalization, first for Artinian local rings, then for local rings whose Jacobson radical are nilpotent, finally as one of the main results of this investigation, we succeeded to show that the same splitting holds for any complete local ring. However, the proof of this general case is technical, and uses the tools and methods of commutative algebra, the theory of fields, and homological algebra. In fact, we have the following result:
@@
The first part of this theorem is well known. In fact, it easily follows from the existence of the coefficient field (see e.g. [8, Theorem 28.3(ii)]). Although the existence of the coefficient fields for equi-characteristic complete local rings is a stronger result, proving their existence is the most difficult part of the proof. However, the second part of the above theorem is a new result.
In our proof there is no need for the existence of the coefficient field. The characteristic of the residue field $k=R/M$ plays an important role in our proof and hence it is necessary to examine it carefully. We know that in any local ring $(R,M,k)$, if $\operatorname{char}(R)=\operatorname{char}(k)$ then there are two cases: $\operatorname{char}(k)=0$ or $\operatorname{char}(k)=p$ for some prime number $p$. However, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then the characteristic of $k$ is always a prime number $p$.
To prove the above theorem, in the equi-characteristic case, one of the main ingredients that applied is the following important result:
@@
To prove Theorem 1.1 in the unequal characteristic case, one of the main tools which used is a key result in homological algebra on the vanishing of Ext groups:
@@ 2. Finite case-the initial motivation
Also recall that if $f:A\rightarrow B$ is a morphism of rings, then it can be seen that $\operatorname{char}(B)$ divides $\operatorname{char}(A)$, i.e., $\operatorname{char}(B)|\operatorname{char}(A)$. In particular, if we have morphisms of rings $A\rightarrow B$ and $B\rightarrow A$, then $\operatorname{char}(A)=\operatorname{char}(B)$.
If we have a ring map $A\rightarrow B$ with $\operatorname{char}(B)=0$, then $\operatorname{char}(A)=0$. However, if $\operatorname{char}(A)=p^{n}$ for some prime number $p$, then $\operatorname{char}(B)=p^{d}$ with $d\leqslant n$.
@@
Recall that by complete local ring we mean a local ring $(R,M)$ such that it is complete with respect to the $M$-adic topology (i.e., $R$ is Hausdorff and every Cauchy sequence is convergent in $R$), or equivalently, the natural ring map $R\rightarrow\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ given by $r\mapsto(r+M^{n})_{n\geqslant 1}$ is an isomorphism. Note that in this definition $R$ is not assumed to be Noetherian (contrary to the classical literature, see e.g. [1], where complete local rings were also assumed to be Noetherian).
@@ 3. The general case
Using the above isomorphism of Abelian groups ${G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle$$\simeq\>\>\>\>\>\>\>\>\>\>\>$${I^{n}/I^{n+1},}$ we can make $G$ into an $R/I$-module (and then this map becomes an isomorphism of $R/I$-modules). Now the statement about the vanishing is clear, because if $M$ is a module over a ring $R$ then $M$ is vanished by $\operatorname{char}(R)$. ∎
@@
If in a local ring $(R,M,k)$ the natural ring map $R\rightarrow k$ admits a splitting (i.e., has a section) then $\operatorname{char}(R)=\operatorname{char}(k)$. We will see that the converse is also true for complete local rings.
@@
${k_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{inc\>\>}$$\scriptstyle{h\>\>\>\>\>\>}$${k\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h^{\prime}}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f\>\>}$${k.}$
@@
The above corollary is quite sufficient for our purpose. However, we should add that this result holds more generally: Let $R$ be a regular Noetherian ring of Krull dimension $n$ and $M,N$ be modules over $R$. If there is an element $r\in R$ such that $r$ is regular on $M$ and $rN=0$, then $\operatorname{Ext}_{R}^{n}(M,N)=0$. In fact, this is proved by the same argument as the above by applying a theorem of Serre [9] which asserts that if $R$ is regular Noetherian of Krull dimension $n$, then for any $R$-modules $M$ and $N$ one has $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>n$.
@@
Now let $p$ be a prime number. Consider the local ring $\mathbb{Z}_{P}$ where $P=p\mathbb{Z}$. By the above, $\mathbb{Z}_{P}/P\mathbb{Z}_{P}\simeq\mathbb{F}_{p}=\mathbb{Z}/p$. Then the natural surjective map between the groups of units $(\mathbb{Z}_{P})^{\ast}\rightarrow\mathbb{F}_{p}^{\ast}$ does not admit a splitting for $p\geqslant 5$, because if this map splits then $\mathbb{F}_{p}^{\ast}$ can be viewed as a subgroup of $(\mathbb{Z}_{P})^{\ast}$, but $\mathbb{Z}_{P}$ does not contain roots of unit of order p1p-1.
But for $p=2,3$ the map $(\mathbb{Z}_{P})^{\ast}\rightarrow\mathbb{F}_{p}^{\ast}$ admits a splitting. We know that for any prime number $p$, the local ring $\mathbb{Z}_{P}$ is not complete. This shows that the above short exact sequence may be split for some incomplete local rings (in other words, the converse of Corollary 3.16 is not true).
@@
First of all, note that if a ring $R$ is complete with respect to the $I$-adic topology for some ideal $I$ of $R$, then it can be seen that $I$ is contained in the Jacobson radical of $R$. In particular, if a ring $R$ is complete with respect to the $M$-adic topology for some maximal ideal $M$ of $R$, then $R$ is a complete local ring, and for complete local rings we proved that the above splittings already exist. Hence, to find the above examples, the ideal $I$ must be a non-maximal ideal.
@@
For the second type of example, for a prime number $p$, consider the ring of $p$-adic integers $\mathbb{Z}_{p}=\lim\limits_{\overleftarrow{n\geqslant 1}}\mathbb{Z}/p^{n}$ which is a complete local ring of mixed characteristic, i.e., $\operatorname{char}(\mathbb{Z}_{p})=0$ but the reside field of $\mathbb{Z}_{p}$ is $\mathbb{F}_{p}$, and take the ideal $I=p^{2}\mathbb{Z}_{p}$ where by $p^{2}$ we mean the element $(p^{2}+p^{n}\mathbb{Z})_{n\geqslant 1}$ of $\mathbb{Z}_{p}$. By Remark 2.5, we have $\mathbb{Z}_{p}/I\simeq\mathbb{Z}/p^{2}$. Then the group of units of $\mathbb{Z}_{p}/I$ is of order p(p1)p(p-1) and by Cauchy theorem has an element of order $p$ (it can be also directly seen that $1+p$ is an element of $(\mathbb{Z}/p^{2})^{\ast}$ of order $p$). But for $p\geqslant 3$ and for each $n\geqslant 1$, by Gauss theorem, $(\mathbb{Z}/p^{n})^{\ast}$ is a cyclic group of order $p^{n-1}(p-1)$ and so there is a natural isomorphism of groups $(\mathbb{Z}/p^{n})^{\ast}\simeq\mathbb{Z}/(p-1)\times\mathbb{Z}/p^{n-1}$, thus the group $(\mathbb{Z}_{p})^{\ast}=\lim\limits_{\overleftarrow{n\geqslant 1}}(\mathbb{Z}/p^{n})^{\ast}\simeq\mathbb{Z}/(p-1)\times(\lim\limits_{\overleftarrow{n\geqslant 1}}\mathbb{Z}/p^{n-1})\simeq\mathbb{F}_{p}^{\ast}\times\mathbb{Z}_{p}$ has no nontrivial $p$-torsion, so there is no splitting in this case.
We say that a local ring $(R,M,k)$ has a coefficient field if there is a ring map $K\rightarrow R$ such that $K$ maps isomorphically to $k$ under the natural ring map $R\rightarrow k$. If a local ring $(R,M,k)$ has a coefficient field $K$, then the natural ring map $f:R\rightarrow k$ admits a splitting. Indeed, we have a ring map $g:K\rightarrow R$ such that $h:=fg$ is an isomorphism, then $1=hh^{-1}=f(gh^{-1})$ and so we have a section $gh^{-1}:k\rightarrow R$.
It is also clear that if a local ring admits a coefficient field then it is equi-characteristic. As mentioned earlier, the converse is also true for complete local rings: Every equi-characteristic complete local ring has a coefficient field. For its proof see e.g. [8, Theorem 28.3(ii)].
Acknowledgments. The author would like to give sincere thanks to Professors Bhargav Bhatt, Brian Conrad, Pierre Deligne, and Melvin Hochster who generously shared their very valuable and excellent ideas with me during the writing of this article.